OpenStudy (anonymous):

If (x+5)^2=0, what is the value of (x-1)(x+1)?

Directrix (directrix):

(x+5)^2=0 --> Look at (x+5). What value of x makes x + 5 = 0? --------- @william0123

OpenStudy (anonymous):

-5

Directrix (directrix):

I messed that up. Just a second.

Directrix (directrix):

Let's knock this out. Cool. -5 it is. Now, look at this: (x-1)(x+1) If x = -5, then what is the value of (x - 1)(x-1)? Hint: ( -5 - 1) * (-5 +1) Take a guess and post.

Directrix (directrix):

What is this equal to: ( -5 - 1) We'll do this by parts.

OpenStudy (anonymous):

Is it 24?

OpenStudy (anonymous):

altogether

Directrix (directrix):

Correct! ( -5 - 1) * (-5 +1) = (-6)*(-4) = 24. Yes, you are correct. Hey, you could multiply (x-1)(x+1) to get x^2 -1 and then substitute -5 and get the same result.

OpenStudy (anonymous):

@Directrix Thanks :)

Directrix (directrix):

Glad to help. I enjoyed it.